OpenStudy (kkutie7):

I have this question... I'm trying to work on it first so you can see my work so it will be up soon. Q: Calculate the % iron in an unknown sample if 0.3400 g unknown was weighed, and 31.45 mL of 0.01788 M K2Cr2O7 was used to titrate to the endpoint. Make sure you use the coefficients from the reaction balanced in problem 2 above.

OpenStudy (kkutie7):

\[6 Fe^{+2} + Cr-{2}O_{7}^{2-} + 14 H+\rightarrow 6 Fe^{+3} + 2 Cr^{+3} + 7 H_{2}O\]

OpenStudy (kkutie7):

my first thought was to do this first \[\frac{.01788 M K_{2}Cr_{2}O_{7}}{1}*\frac{1}{.031450L}=.568521M\]

OpenStudy (kkutie7):

or should I be trying to get moles instead?

OpenStudy (australopithecus):

can you please rewrite the reaction formula? You dont really need to How to solve: 1. Use this info to find moles 31.45mL of 0.01788 M K2Cr2O7, using the formula: Molarity = moles/Volume Volumes has to be in liters so multiple 31.45mL by (1L/1000mL) to convert to liters 2. Use your reaction formula, to find moles of iron 3. convert moles to grams using Moles = grams/molecular mass 4. Use grams of iron found in step 3 and 0.3400g (the weight of the unknown) in the following formula: % iron in unknown = (Grams of Iron/Grams of unknown)*100

OpenStudy (kkutie7):

\[\frac{.01788M}{1L}*.031450L=5.626260*10^{-4}moles\]

OpenStudy (australopithecus):

Can you take a screen shot of the formula or rewrite it because what you provided is pretty illegible

OpenStudy (kkutie7):

really how so it looks perfectly fine for me.

OpenStudy (kkutie7):

\[6 Fe^{+2} + Cr_{2}O_{7}^{2-} + 14 H+\rightarrow 6 Fe^{+3} + 2 Cr^{+3} + 7 H_{2}O\] better?

OpenStudy (australopithecus):

It looks like a mess to me are you sure that is exactly how it is written in your homework. I provided the steps to solve your problem above.

OpenStudy (australopithecus):

Is that exactly how it is written?

OpenStudy (kkutie7):

I'm confused because there is no K2Cr2O7 in the equation. yes that is exactly how it is written.

OpenStudy (australopithecus):

You just need to use gravimetric factor to find moles of Cr2O^2 in K2Cr2O7

OpenStudy (australopithecus):

I will edit step 1 1. a) Use this info to find moles 31.45mL of 0.01788 M K2Cr2O7, using the formula: Molarity = moles/Volume Volumes has to be in liters so multiple 31.45mL by (1L/1000mL) to convert to liters b) use gravimetric factor with the following equation using moles of K2Cr2O7 found above. \[(Moles of K2Cr2O7)*\frac{molecular\; mass\; of\; Cr_2O_7^{-2}}{Molecular\; mass\; of\; K_2Cr_2O_7} = moles\; of\; Cr_2O_7^{-2}\]

OpenStudy (australopithecus):

In step 2. All you have to do is divide moles of Cr2O7^-2 by 1 and then multiply them by 6 to get moles of Fe^2+ then convert to grams of Iron using: moles = grams/molecular mass

OpenStudy (australopithecus):

If you dont know how to find molecular mass look up how to find molecular mass in youtube

OpenStudy (australopithecus):

That is all the information you need to solve this problem

OpenStudy (kkutie7):

\[5.626260*10^{-4}moles of K_2Cr_2O_7^{2-}*\frac{215.985000}{294.181000}=4.130749*10^{-4}molesCr_{2}O_{7}^{2-}\]

OpenStudy (kkutie7):

\[4.130749*10^{-4}molesCr_{2}O_{7}^{2-}*\frac{6molesFe^{2+}}{1molesCr_{2}O_{7}^{2-}}\] \[=.002478molesFe^{2+}\] \[.002478molesFe^{2+}*\frac{55.845gFe^{2+}}{1molesFe^{2+}}\] \[.138384g Fe^{2+}\]

OpenStudy (kkutie7):

\[\frac{.3400}{.138384}*100=\] this isn't right @Australopithecus i didn't work

OpenStudy (australopithecus):

0.01788*(31.45/1000) = 0.000562326mol http://www.wolframalpha.com/input/?i=0.01788*%2831.45%2F1000%29 0.000562326mole*(215.988/294.185) = 0.00041285mol http://www.wolframalpha.com/input/?i=+0.000562326*%28215.988%2F294.185%29 0.00041285mol*6 = 0.0024771mol http://www.wolframalpha.com/input/?i=0.00041285mol*6 0.0024771mol*55.845g/mol = 0.13833g http://www.wolframalpha.com/input/?i=0.0024771mol*55.845g%2Fmol (0.13833/0.138384)*100 = 99.96% http://www.wolframalpha.com/input/?i=+%280.13833%2F0.138384%29*100

OpenStudy (australopithecus):

Molecular mass of Cr2O7 is http://www.wolframalpha.com/input/?i=Cr2O7+molecular+mass Molecular mass of K2Cr2O7 is http://www.wolframalpha.com/input/?i=K2Cr2O7+molecular+mass 294.185 daltons = 294.185g/mol

OpenStudy (australopithecus):

My method is correct I think you just made a mistake in the calculation

OpenStudy (kkutie7):

im with you up till here: 0.13833/0.138384)*100 = 99.96% where did you get the .138384 from?

OpenStudy (kkutie7):

@Australopithecus

OpenStudy (kkutie7):

should it be .3400?

OpenStudy (kkutie7):

I got 41% that way

OpenStudy (australopithecus):

oh yeah you are right

OpenStudy (australopithecus):

0.3400 g is the mass of the unknown my mistake

OpenStudy (australopithecus):

(0.13833g/0.3400g)*100 = 40.68% http://www.wolframalpha.com/input/?i=%280.13833%2F0.3400%29*100+ You are right I was looking at your solution for some reason even though I knew it was wrong. :S

OpenStudy (australopithecus):

that is where I got 0.138384

OpenStudy (australopithecus):

Also I should have noticed you had the ratio flipped the wrong way

OpenStudy (australopithecus):

Dont get confused by my mistakes just follow my steps, if you do it exactly as i have written out you wont have any problems, but pretty sure the answer is 40%